Given the following demand function: Q=120 - 4P + 0.01M, where M is income and stands at $40,000 and the prevailing market price is $10. Calculate the consumer surplus in this market?Given the following demand function: Q=120 - 4P + 0.01M, where M is income and stands at $40,000 and the prevailing market price is $10. Calculate the consumer surplus in this market? What is the market's willingness to pay
1
Expert's answer
2020-09-28T09:50:18-0400

SOLUTION

GIVEN DATA

demand function is

Q=120-4P+0.01M


M is 40000$

so demand function is

Q=520-4P

and P=10 so

Q=520 - 40

Q=480

to see the choke price


Q=0

so

0=520 - 4P

then

P=130


there we have a triangle with height=130-10=120

and base= 480

so

consumer surplus ="\frac{1}{2}\times height\times base"


"consumer\space surpluse=\frac{1}{2}\times120\times480 =28800"